# RA-RLHF: Provably Efficient Risk-Aware Reinforcement Learning Human Feedback

### v2

@@ Abstract
Reinforcement Learning Human Feedback (RLHF) studies the problem where agents receive only preferences over pairs of trajectories in each episode. Traditional approaches in this field have predominantly focused on the mean reward or utility criterion. However, in RLHF scenarios demanding heightened risk awareness, such as in AI systems, healthcare, and agriculture, risk-aware measures are requisite. Traditional risk-aware objectives and algorithms are not applicable in such one-episode-reward settings. To address this, we explore and prove the applicability of two risk-aware objectives to RLHF: nested and static quantile risk objectives. We also introduce Risk-Aware-RLHF (RA-RLHF), an algorithm designed to optimize both nested and static objectives. Additionally, we provide a theoretical analysis of the regret upper bounds, demonstrating that they are sublinear with respect to the number of episodes, and present empirical results to support our findings. Our code is available in https://github.com/aguilarjose11/pbrlNeurips.
@@ 1 INTRODUCTION
Reinforcement Learning (RL) (Russell & Norvig, 2010) is a fundamental framework for sequential decision-making, enabling intelligent agents to interact with and learn from unknown environments. This framework utilizes a reward signal to guide the selection of policies, where optimal policies maximize this signal. RL has demonstrated advanced performance in various domains, including clinical trials (Coronato et al., 2020), gaming (Silver et al., 2017), and autonomous driving (Basu et al., 2017).
@@
Despite its performance, a significant limitation of the standard RL paradigm is the selection of a state-action reward function. In many real-world scenarios, constructing an explicit reward function is often a complex or unfeasible task. As a compelling alternative, Reinforcement Learning Human Feedback (RLHF) (Busa-Fekete et al., 2014; Wirth et al., 2016) addresses this challenge by deviating from traditional quantifiable rewards for each step. Instead, RLHF employs binary preference feedback on trajectory pairs generated by two policies, which can be provided directly by human subjects. This method is increasingly recognized as a more intuitive and direct approach in fields involving human interaction and assessment, such as autonomous driving (Basu et al., 2017), healthcare (Coronato et al., 2020), and language models (Bai et al., 2022).
@@
Previous approaches to RLHF (Xu et al., 2020a; Coronato et al., 2020; Xu et al., 2020b; Chen et al., 2022; Zhan et al., 2023) mainly aim to maximize the mean reward or utility, which is risk-neutral. However, there is a growing need for risk-aware strategies in various fields where RLHF has shown empirical success. For example, in autonomous driving, RLHF reduces the computational burden by skipping the need to calculate reward signals for every state-action pair (Chen et al., 2022). Despite this improvement, the nature of the problem makes dangerous actions costly. Thus, such risk-sensitive problem settings require risk awareness to ensure safety.
@@
Risk-Aware RLHF also has implications for fields like generative AI (OpenAI, 2023; Chen et al., 2023), where harmful content generation remains a challenge for fine-tuning. In this scenario, a large language model (LLM) is often fine-tuned with user feedback by generating two prompt responses, A and B. Many approaches using RLHF (Ouyang et al., 2022) consider human feedback to penalize harmful content generation. Unfortunately, current approaches only minimize the average harmfulness of a response. This can be a challenge when responses are only harmful to a minority of users. Risk-aware RLHF tackles this challenge by directly aiming to decrease the harmfulness directly, rather than indirectly as with human feedback fine-tuning.
@@
Despite substantial evidence highlighting the importance of risk awareness in RLHF, a significant gap persists in the theoretical analysis and formal substantiation of risk-aware RLHF approaches. This deficiency has spurred us to develop risk-aware measures and their theoretical analysis within RLHF. In standard RL, a variety of risk-aware measures have been explored, including a general family of risk-aware utility functions (Fei et al., 2020), iterated (nested) Conditional Value at Risk (CVaR) (Du et al., 2022), and risk-sensitive with quantile function form (Bastani et al., 2022). In general, these measures can be categorized into two types: nested or static. Nested measures (Fei et al., 2020; Du et al., 2022) utilize MDPs to ensure risk sensitivity of the value iteration at each step under the current state, resulting in a more conservative approach. In contrast, static risk-aware measures Bastani et al., 2022 analyze the risk sensitivity of the whole trajectory’s reward distribution. In developing and introducing risk-aware objectives in RLHF, we have encountered the following technical challenges in algorithm design and theoretical analysis:
@@
In RLHF, the reward function depends on the preference between two trajectories generated by the agent. We refer to this difference in how the reward function is computed as the one-episode-feedback characteristic. Consequently, the risk-aware objectives of standard RL like Du et al. (2022) and Fei et al. (2020) become unmeasurable since they depend on the state-action reward.
@@
When using a quantile function to transform a risk-neutral RLHF algorithm into a risk-aware algorithm, the Bellman equation used to solve the problem becomes non-linear. This change to the bellman equation disrupts calculations on regret, making risk-neutral RLHF inapplicable. This is primarily due to the additional parameter $\alpha$, which modifies the underlying distribution.
In this paper, we address these challenges by studying the feasibility of risk-aware objectives in RLHF. We propose a provably efficient algorithm, RA-RLHF, with theoretical and empirical results on its performance and risk-awareness. Our summary of contributions is as follows:
@@
    We analyze the feasibility of several risk-aware measures in RLHF settings and prove that in the one-episode-reward setting, nested and static quantile risk-aware objectives are applicable since they can be solved and computed uniquely in a given RLHF MDP.
@@
    We expand the state space in our formulation of a RLHF MDP and modify value iteration to address its history-dependent characteristics from the one-episode setting. These modifications enable us to use techniques like DPP to search for the optimal policy.
@@
    We develop a provably efficient (both computationally and statistically) algorithm, RA-RLHF, for nested and static quantile risk-aware objectives. To the best of our knowledge, we are the first to formulate and analyze the finite time regret guarantee for a risk-aware algorithm with non-Markovian reward models for both nested and static risk-aware objectives. Moreover, we construct hard-to-learn instances to establish a lower bound for regret.
@@ 2.1 Preference-based Feedback Reinforcement Learning
The incorporation of human preferences in RL, such as Jain et al. (2013), has been a subject of study for over a decade. This approach has proved to be successful and has been widely used in various applications, including language model training (Ouyang et al., 2022), clinical trials (Coronato et al., 2020), gaming (Silver et al., 2017), and autonomous driving (Basu et al., 2017). RLHF can be categorized into three distinct types Wirth et al. (2017): action preference, policy preference, and trajectory preference. Among these, trajectory preference is identified as the most general and widely studied form of preference-based feedback, as evidenced by the rich literature on the topic (Chen et al., 2022; Xu et al., 2020a; Wu & Sun, 2023). As noted in our introduction, previous theoretical explorations on RLHF have predominantly aimed at achieving higher average rewards, which encompasses risk-neutral RLHF. We distinguish our work by taking the novel approach of formalizing the risk-aware RLHF problem.
@@ 2.2 Risk-aware Reinforcement Learning
Practical and relevant trajectory or state-action embeddings are described in works such as Pacchiano et al. (2021). Therefore, the one-episode-reward might not even be sum-decomposable (the trajectory embedding details can be seen in sec.3.1 ). Compared to previous work, we use non-Markovian reward models that do not require estimating the reward at each step and explore both nested and static risk-aware objectives, aiming to provide a more general method.
@@ 3.1 RLHF MDP
Intuitively, a trajectory encapsulates the interactions between an agent and the environment $\mathcal{M}$ up to step $h$. In contrast to the standard RL setting, where the reward function $r^{\star}_{h}(s_{h},a_{h})$ specifies the reward at each step $h$, a significant distinction in the RLHF MDP framework is that the reward function $r^{\star}$ is defined as $r_{\xi}^{\star}(\xi_{H}):(\mathcal{S}\times\mathcal{A})^{H-1}\times\mathcal{S}\rightarrow[0,1]$, denoting the reward of the entire trajectory.
@@
At each iteration $k\in[K]$, the agent selects two policies under a deterministic policy framework, $\pi_{1,k}$ and $\pi_{2,k}$, which generate two (randomized) trajectories $\xi_{H}^{1,k}$ and $\xi_{H}^{2,k}$. In RLHF, unlike standard RL where the agent receives rewards every step, the agent can only obtain the preference $o_{k}$ between two trajectories $\left(\xi_{H}^{1,k},\xi_{H}^{2,k}\right)$. By making a query, we obtain a preference feedback $o_{k}\in\{0,1\}$ that is sampled from a Bernoulli distribution:
@@ 3.2 Risk Measure
Because in RLHF we can only estimate the reward for the entire trajectory, the risk measure selected for RLHF must rely solely on the reward of the entire trajectory. That is, two trajectories with the same trajectory reward should contribute equally to the risk measure, even if their potential rewards at each step are different. Unlike Chen et al. (2023) that decomposes the reward at each step (where the solution is likely not unique) and then calculates the risk measure, this requirement ensures that the risk measure consistently and holistically reflects the underlying preference. We refer to risk measures that suitable for RLHF problems as RLHF-risk-measures. Here, we introduce two different risk-measures: nested and static quantile risk-aware measures, which are appropriate for RLHF-MDPs.
@@
Nested RLHF-risk-measures. For standard RL’s MDP, the nested quantile risk-aware measure can be elucidated in Bellman equation type as follows:
@@
For the RLHF-MDP setting $\mathcal{M}(\mathcal{S},\mathcal{A},r_{\xi}^{\star},\mathbf{P}^{\star},H)$ , the state-action’s reward might not be calculated or the reward of the entire trajectory might not be decomposed. Therefore, the policy should be history-dependent. We rewrite the nested quantile objective’s Bellman equation with the embedded trajectory reward as follows:
@@
For any RLHF MDP $\mathcal{M}(\mathcal{S},\mathcal{A},r_{\xi},\mathbf{P},H)$, we use $\tilde{V}_{h}^{\pi,r_{\xi},\mathbf{P}}(\xi_{h})$ to denote the value iteration under the policy $\pi$, where $\pi$ is a history dependent policy.
@@
Static RLHF-risk-measures.
@@
Also, we modify the distributional Bellman equation for RLHF MDP $\mathcal{M}(\mathcal{S},\mathcal{A},r_{\xi}^{\star},\mathbf{P}^{\star},H)$ settings as follows:
@@ 3.3 Objective
Regardless of nested or static CVaR objectives, we are given an algorithm for computing $\pi_{\mathcal{M}}^{\star}$ for a known $\operatorname{RLHF-MDP}\mathcal{M}$.
@@
4 Risk Aware Reinforcement Learning Human Feedback Algorithm
In this section, we introduce and analyze an algorithm called RA-RLHF for solving the RLHF problem with both nested and static risk aware objectives. Also, we establish a regret bound for it.
@@ 4.1 Algorithm
RA-RLHF is formally described in Algorithm 1. The development of RA-RLHF is primarily inspired by the PbOP algorithm, as delineated in Chen et al. (2022), which was originally proposed for risk-neutral RLHF environments. Building upon this foundation, one significant difference is how to choose a risk aware policy in estimated RLHF MDP, where the value iteration is different. We also use novel techniques to estimate the confidence set and explore for a policy, instead of using a bonus (Chen et al., 2022) (which is difficult to calculate in risk-aware problems) as in standard RL.
Algorithm 1 RA-RLHF
@@ 4.2 Analysis
With at least probability $1-\delta$, the nested quantile risk aware object regret of RA-RLHF is bounded by:
@@
The static quantile risk aware object regret of RA-RLHF is bounded by:
@@
The proof of this theorem is provided in Appendix D.21. Notice that the regret for both the nested risk objective and the static risk objective of Algorithm 1 are sublinear with respect to $K$, making RA-RLHF the first provably efficient algorithm with one-episode-reward for these two objectives. Additionally, compared to Chen et al. (2023), we achieve the goal of having both policies gradually approach optimality. Moreover, in comparison to the nested risk-aware objective, the static objective focuses on the risk measure of the entire distribution, primarily influenced by the Lipschitz coefficient $L_{G}$ of the quantile function and is less constrained by certain specific cases.
@@
The nested quantile risk aware object regret of RA-RLHF is bounded by:
@@
The static quantile risk aware object regret of RA-RLHF is bounded by:

### v3

# RA-PbRL: Provably Efficient Risk-Aware Preference-Based Reinforcement Learning

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has recently surged in popularity, particularly for aligning large language models and other AI systems with human intentions. At its core, RLHF can be viewed as a specialized instance of Preference-based Reinforcement Learning (PbRL), where the preferences specifically originate from human judgments rather than arbitrary evaluators. Despite this connection, most existing approaches in both RLHF and PbRL primarily focus on optimizing a mean reward objective, neglecting scenarios that necessitate risk-awareness, such as AI safety, healthcare, and autonomous driving. These scenarios often operate under a one-episode-reward setting, which makes conventional risk-sensitive objectives inapplicable. To address this, we explore and prove the applicability of two risk-aware objectives to PbRL: nested and static quantile risk objectives. We also introduce Risk-Aware-PbRL (RA-PbRL), an algorithm designed to optimize both nested and static objectives. Additionally, we provide a theoretical analysis of the regret upper bounds, demonstrating that they are sublinear with respect to the number of episodes, and present empirical results to support our findings. Our code is available in https://github.com/aguilarjose11/PbRLNeurips.
@@ 1 INTRODUCTION
Reinforcement Learning (RL) (Russell & Norvig, 2010) is a fundamental framework for sequential decision-making, enabling intelligent agents to interact with and learn from unknown environments. This framework utilizes a reward signal to guide the selection of policies, where optimal policies maximize this signal. RL has demonstrated state-of-the-art performance in various domains, including clinical trials (Coronato et al., 2020), gaming (Silver et al., 2017), and autonomous driving (Basu et al., 2017).
@@
Despite its performance, a significant limitation of the standard RL paradigm is the selection of a state-action reward function. In many real-world scenarios, constructing an explicit reward function is often a complex or unfeasible task. As a compelling alternative, Preference-based Reinforcement Learning (PbRL) (Busa-Fekete et al., 2014; Wirth et al., 2016) addresses this challenge by deviating from traditional quantifiable rewards for each step. Instead, PbRL employs binary preference feedback on trajectory pairs generated by two policies, which can be provided directly by human subjects. This method is increasingly recognized as a more intuitive and direct approach in fields involving human interaction and assessment, such as autonomous driving (Basu et al., 2017), healthcare (Coronato et al., 2020), and language models (Bai et al., 2022).A specialized and increasingly popular variant of PbRL is Reinforcement Learning from Human Feedback (RLHF)Bai et al. (2022), which has garnered considerable attention for aligning AI systems with human intentions—particularly in the domain of large-scale language models.
@@
Previous approaches to PbRL (Xu et al., 2020a; Coronato et al., 2020; Xu et al., 2020b; Chen et al., 2022; Zhan et al., 2023) mainly aim to maximize the mean reward or utility, which is risk-neutral. However, there is a growing need for risk-aware strategies in various fields where PbRL has shown empirical success. For example, in autonomous driving, PbRL reduces the computational burden by skipping the need to calculate reward signals for every state-action pair (Chen et al., 2022). Despite this improvement, the nature of the problem makes dangerous actions costly. Thus, such risk-sensitive problem settings require risk awareness to ensure safety.
@@
Risk-Aware PbRL also has implications for fields like generative AI (OpenAI, 2023; Chen et al., 2023), where harmful content generation remains a challenge for fine-tuning. In this scenario, a large language model (LLM) is often fine-tuned with user feedback by generating two prompt responses, A and B. Many approaches using RLHF (Ouyang et al., 2022) consider human feedback to penalize harmful content generation. Unfortunately, current approaches only minimize the average harmfulness of a response. This can be a challenge when responses are only harmful to a minority of users. Risk-aware PbRL tackles this challenge by directly aiming to decrease the harmfulness directly, rather than indirectly as with human feedback fine-tuning.
@@
Despite substantial evidence highlighting the importance of risk awareness in PbRL, a significant gap persists in the theoretical analysis and formal substantiation of risk-aware PbRL approaches. This deficiency has spurred us to develop risk-aware measures and their theoretical analysis within PbRL. In standard RL, a variety of risk-aware measures have been explored, including a general family of risk-aware utility functions (Fei et al., 2020), iterated (nested) Conditional Value at Risk (CVaR) (Du et al., 2022), and risk-sensitive with quantile function form (Bastani et al., 2022). In general, these measures can be categorized into two types: nested or static. Nested measures (Fei et al., 2020; Du et al., 2022) utilize MDPs to ensure risk sensitivity of the value iteration at each step under the current state, resulting in a more conservative approach. In contrast, static risk-aware measures Bastani et al., 2022 analyze the risk sensitivity of the whole trajectory’s reward distribution. In developing and introducing risk-aware objectives in PbRL, we have encountered the following technical challenges in algorithm design and theoretical analysis:
@@
In PbRL, the reward function depends on the preference between two trajectories generated by the agent. We refer to this difference in how the reward function is computed as the one-episode-feedback characteristic. Consequently, the risk-aware objectives of standard RL like Du et al. (2022) and Fei et al. (2020) become unmeasurable since they depend on the state-action reward.
@@
When using a quantile function to transform a risk-neutral PbRL algorithm into a risk-aware algorithm, the Bellman equation used to solve the problem becomes non-linear. This change to the bellman equation disrupts calculations on regret, making risk-neutral PbRL inapplicable. This is primarily due to the additional parameter $\alpha$, which modifies the underlying distribution.
In this paper, we address these challenges by studying the feasibility of risk-aware objectives in PbRL. We propose a provably efficient algorithm, Risk-Aware-PbRL(RA-PbRL), with theoretical and empirical results on its performance and risk-awareness. Our summary of contributions is as follows:
@@
    We analyze the feasibility of several risk-aware measures in PbRL settings and prove that in the one-episode-reward setting, nested and static quantile risk-aware objectives are applicable since they can be solved and computed uniquely in a given PbRL MDP.
@@
    We expand the state space in our formulation of a PbRL MDP and modify value iteration to address its history-dependent characteristics from the one-episode setting. These modifications enable us to use techniques like DPP to search for the optimal policy.
@@
    We develop a provably efficient (both computationally and statistically) algorithm, RA-PbRL, for nested and static quantile risk-aware objectives. To the best of our knowledge, we are the first to formulate and analyze the finite time regret guarantee for a risk-aware algorithm with non-Markovian reward models for both nested and static risk-aware objectives. Moreover, we construct a hard-to-learn instance for RA-PbRL to establish a regret lower bound.
@@ 2.1 Preference-based Feedback Reinforcement Learning
The incorporation of human preferences in RL, such as Jain et al. (2013), has been a subject of study for over a decade. This approach has proved to be successful and has been widely used in various applications, including language model training (Ouyang et al., 2022), clinical trials (Coronato et al., 2020), gaming (Silver et al., 2017), and autonomous driving (Basu et al., 2017). PbRL can be categorized into three distinct types Wirth et al. (2017): action preference, policy preference, and trajectory preference. Among these, trajectory preference is identified as the most general and widely studied form of preference-based feedback, as evidenced by the rich literature on the topic Chen et al. (2022); Xu et al. (2020a); Wu & Sun (2023). As noted in our introduction, previous theoretical explorations on PbRL have predominantly aimed at achieving higher average rewards, which encompasses risk-neutral PbRL. We distinguish our work by taking the novel approach of formalizing the risk-aware PbRL problem.
@@ 2.2 Risk-aware Reinforcement Learning
Practical and relevant trajectory or state-action embeddings are described in works such as (Pacchiano et al., 2021). Therefore, the one-episode-reward might not even be sum-decomposable (the trajectory embedding details can be seen in sec.3.1 ). Compared to previous work, we use non-Markovian reward models that do not require estimating the reward at each step and explore both nested and static risk-aware objectives, aiming to provide a more general method.
@@ 3.1 PbRL MDP
Intuitively, a trajectory encapsulates the interactions between an agent and the environment $\mathcal{M}$ up to step $h$. In contrast to the standard RL setting, where the reward function $r^{\star}_{h}(s_{h},a_{h})$ specifies the reward at each step $h$, a significant distinction in the PbRL MDP framework is that the reward function $r^{\star}$ is defined as $r_{\xi}^{\star}(\xi_{H}):(\mathcal{S}\times\mathcal{A})^{H-1}\times\mathcal{S}\rightarrow[0,1]$, denoting the reward of the entire trajectory.
@@
At each iteration $k\in[K]$, the agent selects two policies under a deterministic policy framework, $\pi_{1,k}$ and $\pi_{2,k}$, which generate two (randomized) trajectories $\xi_{H}^{1,k}$ and $\xi_{H}^{2,k}$. In PbRL, unlike standard RL where the agent receives rewards every step, the agent can only obtain the preference $o_{k}$ between two trajectories $\left(\xi_{H}^{1,k},\xi_{H}^{2,k}\right)$. By making a query, we obtain a preference feedback $o_{k}\in\{0,1\}$ that is sampled from a Bernoulli distribution:
@@ 3.2 Risk Measure
Because in PbRL we can only estimate the reward for the entire trajectory, the risk measure selected for PbRL must rely solely on the reward of the entire trajectory. That is, two trajectories with the same trajectory reward should contribute equally to the risk measure, even if their potential rewards at each step are different. Unlike Chen et al. (2023) that decomposes the reward at each step (where the solution is likely not unique) and then calculates the risk measure, this requirement ensures that the risk measure consistently and holistically reflects the underlying preference. We refer to risk measures that suitable for PbRL problems as PbRL-risk-measures. Here, we introduce two different risk-measures: nested and static quantile risk-aware measures, which are appropriate for PbRL-MDPs.
@@
Nested PbRL-risk-measures. For standard RL’s MDP, the nested quantile risk-aware measure can be elucidated in Bellman equation type as follows:
@@
For the PbRL-MDP setting $\mathcal{M}(\mathcal{S},\mathcal{A},r_{\xi}^{\star},\mathbf{P}^{\star},H)$ , the state-action’s reward might not be calculated or the reward of the entire trajectory might not be decomposed. Therefore, the policy should be history-dependent. We rewrite the nested quantile objective’s Bellman equation with the embedded trajectory reward as follows:
@@
For any PbRL MDP $\mathcal{M}(\mathcal{S},\mathcal{A},r_{\xi},\mathbf{P},H)$, we use $\tilde{V}_{h}^{\pi,r_{\xi},\mathbf{P}}(\xi_{h})$ to denote the value iteration under the policy $\pi$, where $\pi$ is a history dependent policy.
@@
Static PbRL-risk-measures.
@@
Also, we modify the distributional Bellman equation for PbRL MDP $\mathcal{M}(\mathcal{S},\mathcal{A},r_{\xi}^{\star},\mathbf{P}^{\star},H)$ settings as follows:
@@ 3.3 Objective
Regardless of nested or static CVaR objectives, we are given an algorithm for computing $\pi_{\mathcal{M}}^{\star}$ for a known $\operatorname{PbRL-MDP}\mathcal{M}$.
@@
4 Risk Aware Preference based RL Algorithm
In this section, we introduce and analyze an algorithm called RA-PbRL for solving the PbRL problem with both nested and static risk aware objectives. Also, we establish a regret bound for it.
@@ 4.1 Algorithm
RA-PbRL is formally described in Algorithm 1. The development of RA-PbRL is primarily inspired by the PbOP algorithm, as delineated in Chen et al. (2022), which was originally proposed for risk-neutral PbRL environments. Building upon this foundation, one significant difference is how to choose a risk aware policy in estimated PbRL MDP, where the value iteration is different. We also use novel techniques to estimate the confidence set and explore for a policy, instead of using a bonus (Chen et al., 2022) (which is difficult to calculate in risk-aware problems) as in standard RL.
Algorithm 1 RA-PbRL
@@ 4.2 Analysis
With at least probability $1-\delta$, the nested quantile risk aware object regret of RA-PBRL is bounded by:
@@
The static quantile risk aware object regret of RA-PBRL is bounded by:
@@
The proof of this theorem is provided in Appendix D.21. Notice that the regret for both the nested risk objective and the static risk objective of Algorithm 1 are sublinear with respect to $K$, making RA-PbRL the first provably efficient algorithm with one-episode-reward for these two objectives. Additionally, compared to Chen et al. (2023), we achieve the goal of having both policies gradually approach optimality. Moreover, in comparison to the nested risk-aware objective, the static objective focuses on the risk measure of the entire distribution, primarily influenced by the Lipschitz coefficient $L_{G}$ of the quantile function and is less constrained by certain specific cases.
@@
The nested quantile risk aware object regret of RA-PBRL is bounded by:
@@
The static quantile risk aware object regret of RA-PBRL is bounded by: